OpenStudy (anonymous):

a certain light truck can go around a flat curve having a radius of 150m with a maximum speed of 32.0m/s^2 . With what maximum speed can it go around a curve having a radius of 75.0m?

sam (.sam.):

Use \[v=\frac{2\pi r}{T}\] \[T=\frac{2\pi r}{v}\] -------------------------------------- For radius 150m, \[T=\frac{2\pi (150)}{32}\] \[T=29.4s\] -------------------------------------- For radius 75m, \[v=\frac{2\pi r}{T}\] \[v=\frac{2\pi (75)}{29.4}\] \[v=?\]

OpenStudy (anonymous):

Hm... this is a centripetal acceleration/force problem. The equation for centripetal acceleration for a particle traveling a circular path is: a=v²/r. Since you're given the values of the equation on the right for first set of conditions, a = 6.83m/s². Solve the equation for v with the second set of conditions to yield v = sqrt(a*r)=sqrt(6.83*75)=22.6m/s. By the way, you wrote that the maximum speed is "32m/s²", when in fact speed is given in m/s.